# Conditional PED-ANOVA: Hyperparameter Importance in Hierarchical & Dynamic Search Spaces

### v1

@@ 4.2. Conditional Local HPI via Within-Regime Variance
5. Numerical Experiments
Here, we first empirically validate condPED-ANOVA on synthetic problems to verify that it properly reflects conditional structure, including parameter activation and regime-dependent domains. We choose a synthetic setting because the expected behavior is clear and easy to diagnose, unlike real-world objectives where the “correct” importance patterns are often ambiguous. Further experimental results, such as runtime comparison, are available in Appendix C.
5.1. Common Experimental Setup
We use a fixed evaluation budget of N=1, [^0] N=1{,}000 and draw each sample independently from a uniform distribution over the search space. We consider three choices of $\gamma\in\{1.0,\,0.75,\,0.5\}$. For each choice of $\gamma$, we sweep $\gamma^{\prime}$ from $0.01$ up to $\gamma-0.01$ in increments of $0.01$, and run experiments for all resulting $(\gamma^{\prime},\gamma)$ quantile pairs. Our condPED-ANOVA implementation is built on top of the PED-ANOVA implementation in Optuna (Akiba et al., 2019). We repeat the HPI computation with 10 different random seeds and report the mean and the standard error across runs. Further details are provided in Section D.1.
@@ 5.2.1. Problem Setting
We first consider a simple synthetic optimization problem with a conditional search space, where the active variable switches depending on a gating parameter. The objective function is defined as:
@@
Our goal is to minimize $f$. Here, only one of $(x,y)$ is active depending on the value of $c$. That is, when $c<0.5$, the parameter $x$ is active and $y$ is inactive; when $c\geq 0.5$, vice versa.
@@
The first setting creates a clear separation between the two branches, whereas the second introduces overlap between the active-variable ranges, making the conditional structure less trivially identifiable from the returned objective values.
@@ 5.2.2. Results and Discussion
We first discuss the disjoint-domain setting (Equation 7) with $\gamma=1.0$ (Figure 1). In this figure, around $\gamma^{\prime}\approx 0.5$, the HPI of the gating variable $c$ is the largest, whereas for a tighter target such as $\gamma^{\prime}\approx 0.1$, the HPI shifts to $x$, and for looser targets such as $\gamma^{\prime}\approx 0.9$, the HPI is dominated by $y$. This behavior is intuitive. When $c<0.5$, the objective equals $x$ and is restricted to a low-valued range (i.e., better performance), while when $c\geq 0.5$, the objective equals $y$ and lies in a high-valued range (i.e., worse performance). Accordingly, whether a sample achieves at least median-level performance is largely determined by which branch is selected, which is reflected by the high importance of $c$ around $\gamma^{\prime}\approx 0.5$. On the other hand, to achieve further better performance (smaller $\gamma^{\prime}$), it is no longer sufficient to select the favorable branch; one must further optimize the active variable itself, which explains the increasing importance of $x$ as $\gamma^{\prime}$ decreases. Moreover, for $\gamma^{\prime}<0.5$, top-$\gamma^{\prime}$ performance requires the $c<0.5$ branch and $y$ to be inactive, which is consistent with the (near) zero importance of $y$. Conversely, for $\gamma^{\prime}>0.5$, whether a sample falls into the top-$\gamma^{\prime}$ region is determined by selecting the $c\geq 0.5$ branch, leading $x$ to be inactive and unimportant.
We next discuss the overlapping-domain setting (Equation 8) with $\gamma=1.0$ (Figure 1). In this setting, the domains of $x$ and $y$ overlap, so the objective values are no longer cleanly separated by the gating variable $c$ alone. Consequently, around $\gamma^{\prime}\approx 0.5$ the HPI of $c$ is no longer close to one, and both $x$ and $y$ also exhibit non-negligible importance. As $\gamma^{\prime}$ decreases, the contribution of $y$ persists up to approximately $\gamma^{\prime}\approx 0.2$, which corresponds to the best values attainable by the $y$-branch (near y=2y=-2). For tighter targets beyond this point, achieving top-$\gamma^{\prime}$ performance requires selecting the $c<0.5$ branch, consistent with the importance of $y$ dropping to (near) zero. Within the $c<0.5$ branch, achieving higher performance increasingly depends on the precise value of $x$, which explains why the HPI of $x$ grows as $\gamma^{\prime}$ becomes smaller.
Also, under both domain settings, the resulting HPI is approximately symmetric for $\gamma^{\prime}<0.5$ and $\gamma^{\prime}>0.5$, reflecting that the objective function is symmetric between the top and bottom regions. However, in the region $\gamma^{\prime}>0.5$, changing $c$ can still improve the objective value; consequently, a small but non-zero importance for $c$ remains even around $\gamma^{\prime}\approx 0.0$, leading to a slight asymmetry. This asymmetry arises from the fact that we minimize the objective, which breaks the top–bottom symmetry.
These results suggest that condPED-ANOVA yields sensible importance values for both the gating variable $c$ and the conditioned variables $x$ and $y$ across target levels $\gamma^{\prime}$. The results for varying $\gamma$ are presented and discussed in Section C.1, where we also observe condPED-ANOVA consistently shows reasonable behavior.
@@ 5.3.1. Problem Setting
We consider a synthetic optimization problem in which the domains of hyperparameters change depending on a gating variable.
The objective is to minimize:
@@
Here, both $x$ and $y$ are always present, but their feasible ranges switch according to $c$.
@@ 5.3.2. Results and Discussion
The results are shown in Figure 1. Similar to the disjoint-domain setting (Equation 7) in the conditional activation experiment (Section 5.2), the gating variable $c$ determines whether a sample achieves top-half performance, which is reflected by the high importance of $c$ around $\gamma^{\prime}\approx 0.5$. On the other hand, unlike the conditional activation case, both $x$ and $y$ are always active and contribute to the objective, so their importance grows toward more extreme targets (smaller or larger $\gamma^{\prime}$). Moreover, $x$ is consistently more important than $y$, reflecting its ability to attain more extreme values (around $\pm 7$).
@@ 5.4. Comparison with Existing Methods
(c) Regime-dependent domains (Equation 9 and Equation 10)
Figure 3. Ablation study results obtained using standard local HPI instead of conditional local HPI we defined in Section 4.2. The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
@@ 5.4.1. Baselines
We compare condPED-ANOVA with existing HPI estimators. Since none of the standard HPI methods are designed to handle conditional hyperparameters in a principled manner, we evaluate them under common naive extensions that enable their application to conditional search spaces.
@@
  • Imputation: For samples in which the target hyperparameter is inactive, we impute its value by a default value (the midpoint of its domain) and compute HPI on the resulting completed dataset.
@@
  • Expansion: We treat the hyperparameter as if it were sampled from a single expanded domain that covers all regime-specific ranges, and compute HPI on this unified representation.
The pseudocodes of these naive extensions are provided in Section D.2. Consequently, for conditional activation, we obtain $4\times 2=8$ baseline variants (four HPI methods combined with two naive handling schemes), whereas for regime-dependent domains, we obtain $4\times 1=4$ variants (four HPI methods combined with the expansion scheme), which we compare against condPED-ANOVA. See Section D.2 for further details on the experimental setup.
5.4.2. Results and Discussion
We first analyze the results on the conditional activation objective (Equation 6) under the disjoint-domain setting (Equation 7). As shown in Figure 2, applying simple filtering to the original PED-ANOVA leads to almost identical HPI values for $x$ and $y$. This behavior is clearly undesirable: in order to achieve better performance (i.e., smaller $\gamma^{\prime}$), the hyperparameter $y$ is inactive, and tuning $y$ cannot affect the objective value at all. This failure arises because filtering discards information about how the presence or absence of the variable relates to the overall performance landscape.
When imputation is applied instead, PED-ANOVA assigns a higher HPI to $x$ in the small-$\gamma^{\prime}$ regime, which is qualitatively reasonable. However, the method still assigns a non-negligible importance to $y$ and it becomes comparable to that of $c$, even though $y$ should be inactive in this region. Moreover, for $\gamma^{\prime}>0.5$, the HPI of $x$ and $y$ exhibit substantial instability. These behaviors are caused by forcibly imputing an artificial value (the midpoint of the domain) for inactive configurations, which distorts the empirical sample distribution and introduces spurious dependencies.
Furthermore, in both filtering and imputation extensions of PED-ANOVA, the hyperparameter $c$, which solely determines whether the objective value belongs to the top-performing half, is not properly identified as important. Around $\gamma^{\prime}\approx 0.5$, all hyperparameters are assigned nearly identical HPI values, indicating a complete failure to capture the dominant role of $c$ in controlling performance.
Similar issues are observed for the other baseline estimators. For f-ANOVA and MDI with filtering, all hyperparameters receive identical importance values, while all the remaining variants still assign identical HPI values to $x$ and $y$.
These results demonstrate that naive extensions are insufficient for adapting existing HPI estimators to conditional search spaces and fail to produce meaningful or interpretable importance estimates. In contrast, condPED-ANOVA consistently yields meaningful importance values that reflect the underlying conditional structure (as shown in Section 5.2). Further discussion on the other objectives and domain settings is deferred to Section C.2, where we observe the baselines fail to yield meaningful HPI in conditional settings.
In this section, we conduct ablation studies to examine the impact of key design choices in condPED-ANOVA, focusing on within-regime variance and regime-level weighting.
5.5.1. Effect of Within-Regime Variance for HPI
In Section 4.2, we defined the conditional local HPI using the within-regime local marginal variance $v_{\gamma,\mathrm{within}}^{(d)}$. We therefore compare the resulting HPI estimates with those obtained using the standard local marginal variance $v_{\gamma}^{(d)}$. The closed-form expression for the standard local marginal variance, computed using the indicator function $b_{\gamma^{\prime}}\coloneqq\mathbf{1}\{x\in\mathcal{X}_{\gamma^{\prime}}\}$, is given as follows, where we can also derive a decomposition of the Pearson divergence into within-regime and inter-regime components (cf. Lemma E.2 in Section E.3):
(11) $v_{\gamma}^{(d)}=\left(\frac{\gamma^{\prime}}{\gamma}\right)^{\!2}\Bigg($ $\underbrace{\!\sum_{i=1}^{\;K^{(d)}}\!\!\frac{\big(\alpha_{i}^{(d)}\big)^{2\!}}{\beta_{i}^{(d)}}D_{\mathrm{PE}}\!\left(p_{\gamma^{\prime},\,i}^{(d)}\middle\|p_{\gamma,\,i}^{(d)}\!\right)}_{\text{within-regime divergence}}$
$\hskip 30.00005pt+\underbrace{D_{\mathrm{PE}}\!\left(\!\bm{\alpha}^{(d)}\middle\|\bm{\beta}^{(d)}\!\right)}_{\text{inter-regime divergence}}\Bigg),$
@@ 5.5.1. Effect of Within-Regime Variance for HPI
Figure 4. Ablation study results obtained by using the naive aggregation scheme (Equation 12) instead of our closed-form expression (Theorem 4.2). The bottom row shows the HPIs before normalization, i.e., $\tilde{v}_{\gamma,\mathrm{within}}^{(d)}$. The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
@@
The results obtained using the standard local marginal variance are shown in Figure 3. In both the conditional-activation setting with disjoint domains (Figure 3) and the regime-dependent domain setting (Figure 3), the HPI of the gating variable $c$ is underestimated around $\gamma^{\prime}\approx 0.5$ and becomes lower than those of $x$ and $y$. This behavior is undesirable, since selecting the appropriate branch is crucial for achieving median-level performance. Moreover, in the conditional-activation setting (Figures 3 and 3), the importance of $x$ or $y$ does not vanish even in the regions where the variable is inactive and is assigned non-zero HPIs, which are even larger than that of $c$ (e.g., $y$ for $\gamma^{\prime}<0.5$ and $x$ for $\gamma^{\prime}>0.5$ in Figure 3, and $y$ for $\gamma^{\prime}\lesssim 0.2$ and $x$ for $\gamma^{\prime}\gtrsim 0.8$ in Figure 3). This happens because the inter-regime variance is primarily induced by the gating variable $c$. When HPI is computed using the standard local marginal variance that includes this inter-regime component, the effect of $c$ leaks into the HPI estimates for $x$ and $y$. As a result, $c$ is suppressed by $x$ and $y$ in the region where $c$ should be most important, and $x$ and $y$ receive spurious non-zero importance even in the regions where they should be inactive and have zero effect. This phenomenon can also be established mathematically. Theorem E.1 shows that, under the standard local marginal variance, the variance of a conditioned hyperparameter includes the gating variance as an additive term, which explains the observed leakage and the resulting misattribution of importance. Therefore, in conditional settings, it is essential to compute HPI using the within-regime local marginal variance $v_{\gamma,\mathrm{within}}^{(d)}$, rather than the standard local marginal variance $v_{\gamma}^{(d)}$.
@@ 5.5.2. Effect of Regime-Level Weighting for HPI
In Section 4.3, we derived a closed-form computation method for $v_{\gamma,\mathrm{within}}^{(d)}$ based on a mathematical derivation (Theorem 4.2). However, one might wonder whether such a derivation is necessary in practice, and instead consider a naive alternative that simply sums the within-regime Pearson divergences, i.e., using:
@@
(12) $\tilde{v}_{\gamma,\mathrm{within}}^{(d)}=\left(\frac{\gamma^{\prime}}{\gamma}\right)^{\!2}\sum_{i=1}^{K^{(d)}}D_{\mathrm{PE}}\!\left(p_{\gamma^{\prime},\,i}^{(d)}\,\middle\|\,p_{\gamma,\,i}^{(d)}\right)$
in place of Equation 5. In this section, we empirically compare condPED-ANOVA against this naive aggregation scheme.
@@
The results obtained using this naive aggregation scheme are shown in Figure 4. In this case, the resulting HPI curves become noticeably less stable and exhibit near-discontinuous changes around regime-switching boundaries. In particular, for the conditional-activation objective under the overlapping-domain setting (Equations 6 and 8), the importance of $y$ spikes immediately when $y$ becomes active, excessively dominating the HPI and suppressing the importances of $x$ and $c$, which should still contribute to the objective. Moreover, although the objective is symmetric between the top region ($\gamma^{\prime}<0.5$) and the bottom region ($\gamma^{\prime}>0.5$) across all settings, the naive scheme breaks this symmetry too severely, yielding clearly distorted importance patterns.
@@
This instability is largely driven by abrupt scale changes at regime boundaries. As shown in the bottom row of Figure 4, $\tilde{v}_{\gamma,\mathrm{within}}^{(d)}$ for $x$ and $y$ exhibit implausible jumps at these boundaries, in sharp contrast to the regime-probability-weighted condPED-ANOVA (cf. Figure 9 in Section C.4). This failure occurs because removing the regime-probability ratios destroys the frequency-aware weighting across regimes, leading to an inappropriate aggregation that cannot properly account for regime prevalence.

### v2

@@ 5. Experiments on Synthetic Problems
Here, we first empirically validate condPED-ANOVA on synthetic problems to verify that it properly reflects conditional structure, including parameter activation and regime-dependent domains. We choose a synthetic setting because the expected behavior is clear and easy to diagnose, unlike real-world objectives where the “correct” importance patterns are often ambiguous. After that, we confirm that condPED-ANOVA also properly works on real-world benchmarks, which is presented in Section 6. Further experimental results, such as runtime comparison and sensitivity analyses with respect to $N$, are available in Appendix D.
5.1. Experimental Setup
We use a fixed evaluation budget of N=1, [^0] N=1{,}000 and draw each sample independently from a uniform distribution over the search space. We consider three choices of $\gamma\in\{1.0,\,0.75,\,0.5\}$. For each choice of $\gamma$, we sweep $\gamma^{\prime}$ from $0.01$ up to $\gamma-0.01$ in increments of $0.01$, and run experiments for all resulting $(\gamma^{\prime},\gamma)$ quantile pairs. Our condPED-ANOVA implementation is built on top of the PED-ANOVA implementation in Optuna (Akiba et al., 2019). We repeat the HPI computation with 10 different random seeds and report the mean and the standard error across runs. To ensure reproducibility, we publicly release our condPED-ANOVA implementation and all experimental codes [^1] https://github.com/kAIto47802/condPED-ANOVA. Further details are provided in Section C.1.
@@ 5.2.1. Problem Setting
We first consider a simple synthetic problem with a conditional search space, where the active variable is selected by a gating parameter. The objective function is defined as:
@@
Our goal is to minimize $f$. When $c<0.5$, the parameter $x$ is active and $y$ is inactive; when $c\geq 0.5$, vice versa.
@@
The first setting cleanly separates the two branches, whereas the second introduces overlap, making the conditional structure less trivially identifiable from the objective values. Experiments with multi-level conditional hierarchies are presented in Section E.1.
@@ 5.2.2. Results and Discussion
We first discuss the disjoint-domain setting (Equation 13) with $\gamma=1.0$ (Figure 1). Around $\gamma^{\prime}\approx 0.5$, the gating variable $c$ has the largest HPI, whereas for a tighter target (e.g., $\gamma^{\prime}\approx 0.1$), the HPI shifts to $x$, and for looser targets (e.g., $\gamma^{\prime}\approx 0.9$), the HPI is dominated by $y$. This behavior is intuitive. When $c<0.5$, the objective equals $x$ and lies in a low (better) range; when $c\geq 0.5$, it equals $y$ and lies in a high (worse) range. Thus, achieving roughly median performance is mainly determined by the selected branch, explaining the high importance of $c$ near $\gamma^{\prime}\approx 0.5$. For smaller $\gamma^{\prime}$, selecting the favorable branch is not enough; one must optimize the active variable, which explains the increasing importance of $x$. Moreover, for $\gamma^{\prime}<0.5$, top-$\gamma^{\prime}$ performance requires the $c<0.5$ branch and $y$ to be inactive, which is consistent with the (near) zero importance of $y$. Conversely, for $\gamma^{\prime}>0.5$, the $c\geq 0.5$ branch dominates, making $y$ important and $x$ inactive (hence unimportant).
We next discuss the overlapping-domain setting (Equation 14) with $\gamma=1.0$ (Figure 1). Because the domains of $x$ and $y$ overlap, objective values are no longer cleanly separated by the gating variable $c$. Consequently, around $\gamma^{\prime}\approx 0.5$ the HPI of $c$ is no longer close to one, and both $x$ and $y$ also have non-negligible HPIs. As $\gamma^{\prime}$ decreases, the contribution of $y$ persists until about $\gamma^{\prime}\approx 0.2$, reflecting the best values attainable in the $y$-branch (near y=2y=-2). For further tighter targets, top-$\gamma^{\prime}$ performance requires the $c<0.5$ branch, consistent with the (near) zero importance of $y$. Within this branch, achieving higher performance increasingly depends on the precise value of $x$, leading to a growing HPI for $x$ as $\gamma^{\prime}$ decreases.
Also, under both domain settings, the resulting HPI is approximately symmetric for $\gamma^{\prime}<0.5$ and $\gamma^{\prime}>0.5$, reflecting the objective’s top–bottom symmetry. However, for $\gamma^{\prime}>0.5$, changing $c$ can still improve the objective, so $c$ retains a non-zero HPI even near $\gamma^{\prime}\approx 1.0$, reflecting a slight asymmetry caused by minimization.
These results suggest that condPED-ANOVA yields sensible HPIs for $c$, $x$, and $y$ across target levels $\gamma^{\prime}$. The results for varying $\gamma$ are presented and discussed in Section D.2, where condPED-ANOVA consistently shows reasonable behavior.
@@ 5.3.1. Problem Setting
We next consider a synthetic optimization problem where hyperparameter domains change with a gating variable. The objective is to minimize:
@@
Here, both $x$ and $y$ are always active, but their domains switch according to $c$. Additional experiments combining conditional activation with regime-dependent domains are provided in Section E.2.
@@ 5.3.2. Results and Discussion
The results are shown in Figure 1. Similar to the disjoint-domain setting (Equation 13) in the conditional activation experiment (Section 5.2), the gating variable $c$ largely determines top-half performance, yielding high importance near $\gamma^{\prime}\approx 0.5$. Unlike that case, however, both $x$ and $y$ are always active and contribute to the objective, consistent with their importance grows toward more extreme targets (smaller or larger $\gamma^{\prime}$). Moreover, $x$ is consistently more important than $y$, reflecting its ability to attain more extreme values (around $\pm 7$).
@@ 5.4.1. Baselines
We then compare condPED-ANOVA with existing HPI estimators. Since these HPI estimators do not handle conditional hyperparameters in a principled way, we evaluate them using common naive extensions for conditional search spaces.
@@
  • Imputation: For samples where the target hyperparameter is inactive, we impute it with a default value (the domain midpoint) and compute HPI on the completed dataset.
@@
  • Expansion: We treat the hyperparameter as if it were sampled from a single expanded domain covering all regime-specific ranges, and compute HPI on this unified domain.
The pseudocodes of these naive extensions are provided in Section C.2. Consequently, for conditional activation we evaluate 8 baseline variants (4 HPI methods $\times$ 2 handling schemes), while for regime-dependent domains we evaluate 4 variants (4 methods $\times$ 1 expansion scheme), and compare them against condPED-ANOVA. See Section C.2 for further details on the experimental setup.
@@
Figure 3. Ablation study results using standard local HPI instead of our conditional local HPI. The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
(c) Regime-dependent domains (Equation 15 and Equation 16)
Figure 4. Ablation study results using the naive aggregation scheme (Equation 18) instead of our expression (Theorem 4.2). The bottom row shows the HPIs before normalization, i.e., $\tilde{v}_{\gamma,\mathrm{within}}^{(d)}$. The lines denote the mean, and the shaded regions denote the standard error, both computed over ten independent runs with different random seeds.
5.4.2. Results and Discussion
We first analyze the results on the conditional activation objective (Equation 12) under the disjoint-domain setting (Equation 13), which are shown in Figure 2. As shown in this figure, applying simple filtering to PED-ANOVA leads to almost identical HPIs for $x$ and $y$. This behavior is clearly undesirable: in order to achieve better performance (i.e., smaller $\gamma^{\prime}$), $y$ is inactive, and tuning $y$ cannot affect the objective at all. This failure arises because filtering obscures how the variable relates to performance over the full search space. When imputation is applied instead, the HPIs of $x$ and $y$ exhibit substantial instability for $\gamma^{\prime}>0.5$. These behaviors are caused by forcibly imputing an artificial value for inactive samples, which distorts the sample distribution and introduces spurious dependencies. Furthermore, under both filtering and imputation, PED-ANOVA fails to identify $c$ as important: around $\gamma^{\prime}\approx 0.5$, it assigns nearly identical HPIs to all hyperparameters, even though $c$ largely determines whether a sample falls in the top-performing half.
Similar issues are observed for the other baselines. For f-ANOVA and MDI with filtering, all hyperparameters receive identical HPIs, while all the remaining variants assign identical HPIs to $x$ and $y$.
These results show that naive extensions are insufficient for applying existing HPI estimators to conditional search spaces, often yielding uninterpretable HPIs. In contrast, condPED-ANOVA consistently yields meaningful HPIs that reflect the underlying conditional structure (as shown in Section 5.2). Additional results for other objectives and domain settings are deferred to Section D.4, where the baselines similarly fail in conditional settings.
In this section, we conduct ablation studies to examine the impact of key design choices in condPED-ANOVA.
5.5.1. Effect of Within-Regime Variance for HPI
In Section 4.2, we defined the conditional local HPI using the within-regime local marginal variance $v_{\gamma,\mathrm{within}}^{(d)}$. We therefore compare the resulting HPI estimates with those obtained using the standard local marginal variance $v_{\gamma}^{(d)}$. The closed-form estimation for the standard local marginal variance, computed using the indicator function $b_{\gamma^{\prime}}\coloneqq\mathbf{1}\{x\in\mathcal{X}_{\gamma^{\prime}}\}$, is given as follows, where we can also derive a decomposition of the Pearson divergence into within-regime and inter-regime components (cf. Lemma F.2 in Section F.3):
(17) $v_{\gamma}^{(d)}\!=\left(\frac{\gamma^{\prime}}{\gamma}\right)^{\!2}\!\Bigg(\underbrace{\!\sum_{i=1}^{\;K^{(d)}}\!\!\frac{\big(\alpha_{i}^{(d)}\big)^{2\!}}{\beta_{i}^{(d)}}D_{\mathrm{PE}}\!\left(p_{\gamma^{\prime},\,i}^{(d)}\middle\|p_{\gamma,\,i}^{(d)}\!\right)}_{\text{within-regime divergence}}+\hskip-3.0pt\underbrace{D_{\mathrm{PE}}\!\left(\bm{\alpha}^{(d)}\middle\|\bm{\beta}^{(d)}\right)}_{\text{inter-regime divergence}}\hskip-3.0pt\Bigg),\!$
The results obtained using the standard local marginal variance are shown in Figure 3. In both the conditional-activation with disjoint domain setting (Figure 3) and the regime-dependent domain setting (Figure 3), the HPI of the gating variable $c$ is underestimated around $\gamma^{\prime}\approx 0.5$ and even falls below those of $x$ and $y$, despite branch selection being crucial for median-level performance. Moreover, in the conditional-activation setting (Figures 3 and 3), $x$ and $y$ retain non-zero HPI even where they are inactive (e.g., $y$ for $\gamma^{\prime}<0.5$ and $x$ for $\gamma^{\prime}>0.5$ in Figure 3).
@@ 5.5.1. Effect of Within-Regime Variance for HPI
This happens because the inter-regime variance is primarily induced by the gating variable $c$. When HPI is computed using the standard local marginal variance that includes this inter-regime component, the effect of $c$ leaks into the HPI of $x$ and $y$, suppressing $c$ and assigning spurious HPIs to $x$ and $y$. In fact, we can theoretically show that, under the standard local marginal variance, the variance of a conditioned hyperparameter includes the gating variance as an additive term, which explains the observed leakage and the resulting misattribution of importance (cf. Theorem F.1). Therefore, in conditional settings, it is essential to compute HPI using the within-regime local marginal variance $v_{\gamma,\mathrm{within}}^{(d)}$ rather than the standard local marginal variance $v_{\gamma}^{(d)}$.
@@ 5.5.2. Effect of Regime-Level Weighting for HPI
In Section 4.3, we derived a closed-form method to compute $v_{\gamma,\mathrm{within}}^{(d)}$ (Theorem 4.2). However, one might wonder whether such a derivation is necessary in practice, and instead consider a naive alternative that simply sums the within-regime Pearson divergences, i.e., using:
@@
(18) $\tilde{v}_{\gamma,\mathrm{within}}^{(d)}\coloneqq\left(\frac{\gamma^{\prime}}{\gamma}\right)^{\!2}\sum_{i=1}^{K^{(d)}}D_{\mathrm{PE}}\!\left(p_{\gamma^{\prime},\,i}^{(d)}\,\middle\|\,p_{\gamma,\,i}^{(d)}\right)$
in place of Equation 11.
Figure 5. condPED-ANOVA HPIs on a real-world problem from YAHPO Gym rbv2_super scenario (instance ID 1053) (Pfisterer et al., 2022; Binder et al., 2020). The task is model selection over ranger, aknn, svm, xgboost, rpart, and glmnet: the hyperparameter learner_id selects the learner, activating only the corresponding learner-specific hyperparameters (prefixed by the learner name). The bars denote the mean, and the error bars denote the standard error, both computed over ten independent runs with different random seeds.
@@
The results obtained using this naive aggregation scheme are shown in Figure 4. The resulting HPI curves become noticeably less stable, with near-discontinuous jumps at regime-switching boundaries. In particular, for the conditional-activation objective under the overlapping-domain setting (Figure 4), the importance of $y$ spikes immediately when $y$ becomes active, excessively dominating the HPI and suppressing $x$ and $c$, which should still contribute to the objective. Moreover, although the objective is symmetric between the top ($\gamma^{\prime}<0.5$) and bottom ($\gamma^{\prime}>0.5$) region across all settings, the naive scheme breaks this symmetry too severely, yielding clearly distorted importance patterns.
@@
This instability is driven by abrupt scale changes at regime boundaries. As shown in the bottom row of Figure 4, $\tilde{v}_{\gamma,\mathrm{within}}^{(d)}$ for $x$ and $y$ exhibit implausible jumps at these boundaries, in contrast to the regime-probability-weighted condPED-ANOVA (cf. Figure 11 in Section D.5). This failure occurs because removing the regime-probability ratios destroys the frequency-aware weighting, leading to an inappropriate aggregation that ignores regime prevalence.
6. Experiments on Real-World Benchmarks
6.1. Experimental Setup
We validate that condPED-ANOVA works properly in real-world settings using YAHPO Gym (Pfisterer et al., 2022). We use the rbv2_super scenario (Binder et al., 2020), as it defines a conditional search space in a single-objective setting. We use the instance ID 1053 from the benchmark suites proposed with YAHPO Gym. This benchmark represents a model selection problem over six learners: aknn (approximate kkNN), glmnet (regularized linear model), ranger (random forest), rpart (decision tree), svm (support vector machine), and xgboost (gradient-boosted trees). A categorical hyperparameter learner_id selects the learner, and only the corresponding learner-specific hyperparameters are active under that choice. We use a sample size of $N=1000$ and maximize the accuracy (acc) as the objective. We use $\gamma=1.0$ and $\gamma^{\prime}=0.1$. Further details of the experimental setup, including details of the YAHPO Gym, are provided in Section C.3.
Table 1. Summary statistics of the objective value (acc) for each learner_id in rbv2_super (instance ID 1053). Values are reported as mean $\pm$ standard error, both computed over 10 independent runs with different random seeds. The learner_ids are sorted from best to worst max performance.
Learner ID Min Mean Max
svm $0.727\pm 0.004$ $0.796\pm 0.001$ $0.926\pm 0.005$
xgboost $0.591\pm 0.011$ $0.765\pm 0.001$ $0.853\pm 0.004$
rpart $0.736\pm 0.001$ $0.774\pm 0.001$ $0.846\pm 0.003$
glmnet $0.731\pm 0.001$ $0.778\pm 0.001$ $0.842\pm 0.003$
ranger $0.738\pm 0.002$ $0.777\pm 0.000$ $0.815\pm 0.002$
aknn $0.174\pm 0.001$ $0.591\pm 0.006$ $0.809\pm 0.002$
6.2. Results and Discussion
The results are shown in Figure 5. We observe that trainsize and learner_id, as well as svm- and xgboost-related hyperparameters, have relatively large HPIs, indicating that these factors are most influential. For reference, we report summary statistics of the objective values (acc) for each selected learner in Table 1. The table shows large differences in attainable performance across learners, making it reasonable for larger HPI on learner_id. Moreover, the ordering of learners that appear in the computed HPIs (via learner-specific hyperparameters) in Figure 5 largely matches the learners’ performance ordering in Table 1 (e.g., svm higher and aknn lower), suggesting that condPED-ANOVA correctly captures each learner’s relative impact on performance. Furthermore, within each learner, condPED-ANOVA assigns high HPIs to hyperparameters that are empirically known to be important, such as xgboost’s booster, nrounds and subsample, and svm’s cost, which is reasonable. These results indicate that condPED-ANOVA can effectively identify important hyperparameters in real-world model selection problems with conditional hyperparameters. Experiments on other problems in YAHPO Gym are reported in Appendix G, where condPED-ANOVA constantly produces reasonable HPIs.